Fermi National Laboratory


Accelerator Update for May 27, 2005 - June 1, 2005

Friday May 27
The day shift began with Operations monitoring store 4176 and the Recycler stash, with the Antiproton Source (Pbar) stacking, and with the MiniBooNE, NuMI, and the Meson MCenter experiments taking beam.

Operations terminated store 4176 at 12:01 PM.

At 12:17 PM a Tevatron (TeV) RF station (TRF8) tripped off. An expert investigated and had the station back on ten minutes later, but he found small LCW leak, which he will repair after the TeV studies end.

TeV studies began at 12:26 PM and ended at 2:09 PM. TRF8 repairs began.

Operations put the TeV into shot setup at 2:40 PM.

Operations established combination store 4178 at 4:46 PM with an initial luminosity of 91.43E30.

At 5:05 PM, Operations reported a low stack rate. Two experts investigated and at 11:17 PM reported finding a problem with a Debuncher RF station (DRF8).

Saturday May 28
The midnight shift began with Operations monitoring store 4176, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

The Recycler began stashing antiprotons at 1:17 AM. After the stashing ended, Pbar resumed stacking at 2:06 AM.

At 7 AM, Operations reported a rise in the Booster tunnel humidity. Operators suspected a LCW leak and turned off the Booster and prepared the tunnel for an access. Operators found the leak in period 2 and valved off the LCW. Operations notified experts.

At 10:01 AM, Operations saw evidence of kicker sparking on a scope. Operators investigated and then called in two Booster experts.

The Booster LCW leak was repaired by 12:42 PM.

The Booster kicker was repaired by 12:54 PM.

Operations reestablished Booster beam at 3:03 PM. Pbar resumed stacking, and MiniBooNE, NuMI, and MCenter all resumed taking beam over the next fifteen minutes.

Operations terminated store 4178 around 10 PM and then put the TeV into shot setup.

Sunday May 29
The midnight shift began with Operations monitoring the stash, with Pbar stacking, with the TeV in shot setup, and with MiniBooNE taking beam.

Operations established combination store 4179 at 1 AM with an initial luminosity of 82.2E30. Pbar resumed stacking and NuMI and MCenter resumed taking beam.

The Recycler began stashing antiprotons at 10:44 AM and finished by 11:11 AM.

Around 8:18 PM, store 4179 aborted. There was no quench. Operators investigated and discovered an open VCB in TeV sector A3. Operators called in an expert who reported at 10:09 PM that he had replaced an AC controller for A3 and closed the breaker.

Operations put the TeV into shot setup at 11:08 PM.

Monday May 30
The midnight shift began with Operations monitoring the stash, with the TeV in shot setup, and with MiniBooNE taking beam.

Operations established combination store 4181 at 3:41 AM with an initial luminosity of 79.45E30. Pbar resumed stacking and NuMI and MCenter resumed taking beam.

The Recycler began stashing antiprotons at 9 AM and finished by 9:38 AM.

At 3:30 PM, Operations reported a Debuncher pulsed magnet (D:PMAG) at AP0 tripped off due to a ground fault. Operators couldn't reset the power supply. They called in experts who said the magnet would have to be replaced. This problem halted stacking.

At 7:17 PM, MiniBooNE experimenters reported that they were having vacuum problems and had to stop taking beam. The Duty Mechanic investigated. Operations resumed sending them beam around 11 PM.

Tuesday May 31
The midnight shift began with Operations monitoring store 4181 and the Recycler stash, with Pbar off due to the pulsed magnet replacement, and with MiniBooNE, NuMI, and MCenter taking beam.

The PMAG replacement was completed around 7:44 AM. Pbar resumed stacking.

Operations terminated store 4181 at 8 AM. TeV experts began a period of studies.

At 8:18 AM, Operations reported that a TeV RF station (TRF5) tripped off. Operations called an expert who investigated and had the station back up by 9:22 AM. TeV experts then began a study period.

Operations established beam to the Meson MTest experiment at 12:44 PM.

Operations put the TeV into shot setup at 5:31 PM.

Operations established combination store 4185 at 9:11 PM with an initial luminosity of 96.0E30. Pbar resumed stacking and NuMI resumed taking beam. Beam to NuMI, MCenter and MTest resumed at 11:14 PM.

Wednesday June 1
The midnight shift began with Operations monitoring store 4185, with Pbar stacking, and with MiniBooNE, NuMI, MCenter, and MTest taking beam.

The Recycler began stashing antiprotons at 4:58 AM and ended around 6 AM. Pbar resumed stacking, and MiniBooNE, NuMI, MCenter, and MTest resumed taking beam.

The Plan for Wednesday
The plan for today is to keep this store in until the midnight shift on Thursday, and to stack and stash until then.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 6/1/2005   email Fermilab